OpenStudy (anonymous):

if f(2)=3 and f'(2)=5 find and equation of the tangent line and the normal line to the graph of y=f(x) I know satellite73 answered this, but I am curious what the normal line's equation is and how to find it

myininaya (myininaya):

Normal line is the line perpendicular to the tangent line and goes through the point at (2,3).

myininaya (myininaya):

Or whatever point of tangency we are talking about (2,3) being the one here.

OpenStudy (anonymous):

Okay, so the inverse of the tangent line: y=5x-7

OpenStudy (anonymous):

What's the easiest way to find the inverse?

myininaya (myininaya):

You want to find the inverse now?

OpenStudy (anonymous):

Yeah because I already have found the tangent line.

myininaya (myininaya):

Since lines are 1 to 1, all you need to do is solve for x. Replace x with f^(-1)(x) and y with x. I think you are confused. Why are we finding the inverse of the line?

OpenStudy (anonymous):

Sorry, I thought that was just another name for the normal line... >.<

myininaya (myininaya):

Nope :p

myininaya (myininaya):

So after we find the slope of the tangent line. Let's call that slope m. Then the slope of the normal line with be -1/m.

myininaya (myininaya):

Lines have form y=slope*x+b

myininaya (myininaya):

The slope of the normal line is -1/m So we have y=-1/m*x+b

OpenStudy (anonymous):

So it's just -1/5x-7?

myininaya (myininaya):

Did you put your point into find the y-intercept?

OpenStudy (anonymous):

o.o

myininaya (myininaya):

y=-1/m*x+b The point this lines goes through or should go through is (2,3) put that point in 3=-1/m*2+b solve for b. 3=-2/m+b 3+2/m=b So you have y=-1/m*x+(3+2/m)

OpenStudy (anonymous):

So, y=-1/5x+17/5 ?

myininaya (myininaya):

yep

OpenStudy (anonymous):

Cool ^w^ Thank you!